Solve by the substitution method 2x+3y=8
-6x+y=36 What is the solution set I came up with ( -2,6) am I close

Answer by Alan3354(30924)

Solve by the substitution method
2x+3y=8
-6x+y=36
What is the solution set I came up with ( -2,6) am I close
-----------
Sub -2 for x and 6 for y, see if they're right.
*************************************************
There's no "close" in these, it's right or wrong.
*************************************************
2x+3y=8
2*(-2) + 3*6 =? 8
8 = 8 That one's good.
-----
-6x+y=36
-6*(-2) + 6 =? 36
12 + 6 <> 36
Not right.
----------
How did you get that?